The transverse displacement of an harmonic wave on a stretched rope is y = 0.05 cos(2.9 x - 5.8 t), where x and y are in meters and t is in seconds.

1)

What is the amplitude of this wave?

A =

m

2)

What is the wavelength of this wave?

l =

m

3)

What is the speed with which this wave travels?

|v| =

m/s

4)

In what direction is this wave propagating?

+x

-x

+y

-y

+z

-z

none of the above

Your submissions:

A

Submitted:

Monday, April 25 at 7:23 PM

Feedback:

Right answer !

5)

What is the frequency (NOT the angular frequency) of this wave?

f =
Answer #1

compare the given equation with

y = A*cos(k*x - w*t)

1) A = 0.05 m

2) k = 2.9

2*pi/lamda = 2.9

lamda = 2*pi/2.9

= 2.17 m

3) v = w/k

= 5.8/2.9

= 2 m/s

4) +x


5) f = w/(2*pi)

= 5.8/(2*pi)

= 0.923 Hz